OpenStudy (anonymous):

The percent increase of an item that jumped from $27 to $30 is approximately _____.

OpenStudy (anonymous):

30-27 is how much then take that number and divide it by 27 and find a percent that is close to that 30-27 = 3 3/27 = 1/9 1/9 is greater than 1/10(10%) but less than 1/8(12.5%) so it is in between 10% and 12.5%, or 9*11% = 99% , 99% is almost 100%